# Taking the Next Step with Generative Artificial Intelligence: The Transformative Role of Multimodal Large Language Models in Science Education

### v2

@@ 4. Challenges and Risks of MLLMs in Science Education
While the adaptability of MLLMs in science education is promising, it is essential to approach their potential with caution. The allure of providing students with extensive flexibility in choosing the modality for their learning environment can be strong. However, evidence suggests that minimal guidance in science education may not always yield optimal learning outcomes (Kirschner et al.,, 2006; SWELLER et al.,, 2007). Learners might require guidance in selecting the most appropriate modality to enhance their educational experience effectively. An abundance of options can not only serve as a distraction but can also elevate the overall cognitive load, potentially hindering the learning process. This is especially true for students with low self-regulation capabilities (Lan,, 1998; Ley and Young,, 1998). Consequently, the roles of the teachers remain pivotal in guiding the learning process (Ley & Young, 2001). They play a critical role in facilitating the effective use of these advanced tools, ensuring that the technology enhances, rather than impedes, the learning experience. In this context, the nuanced integration of MLLMs calls for a balanced approach, where the technology is leveraged thoughtfully to complement and augment traditional educational practices rather than replace them outright.
@@
Students should engage in scientific practices in problem-based learning and open inquiries. But minimal guidance can hinder effective learning in science education (Kirschner et al.,, 2006; SWELLER et al.,, 2007). MLLMs should be designed to be adjusted by the teacher to the individual student’s competency level. Therefore, they are able to offer degrees of guidance or scaffolding rather than acting as a personalized, adaptive, and multimodal agent that provides immediate answers and solutions. This mediated approach could encourage students to engage in critical thinking and problem-solving, fostering a deeper understanding and promoting self-directed learning without overexerting the student. The MLLMs should be aligned with educational and instructional strategies that prioritize the development of students’ analytical skills in tandem with their knowledge acquisition.
MLLMs sometimes still generate biased (Venkit et al.,, 2023)incorrect or faulty content, particularly in the context of logical and mathematical operations (Imani et al., 2023b, ). It’s crucial for teachers to be aware of these limitations. Moreover, if students are encouraged to use MLLMs independently, it’s essential that teachers communicate these flaws to them.
@@
With AI becoming an increasingly integral part of our daily lives, policymakers have begun to implement policies to regulate and safeguard AI development and use (Schiff et al.,, 2020). E.g., the currently presented so-called European AI Act ((European Union,, 2023)) requires generative AI systems, such as MLLMs, to disclose when content is AI-generated, design models to prevent the generation of illegal content, and publish summaries of copyrighted data used for training. The use of such AI systems in education will be classified as high-risk applications and will require registration in an EU database.
Ethical considerations encompass the moral and ethical implications of using MLLMs, including issues of bias, privacy, consent, and the overall impact on the quality and equity of education. Fostering a culture of ethical data handling among educational stakeholders, alongside creating awareness about the potential risks and safeguards among the student populace, is crucial to ensure a secure and responsible utilization of MLLMs in educational settings. On every level of deploying MLLMs, developers, educators, administrators, and policymakers, one has to make sure these ethical considerations have to be a priority.
@@
Teachers and learners have a variety of pre- and misconceptions, fears, and hopes about AI (Bewersdorff et al.,, 2023), which might lead to general skepticism toward AI systems in the classroom among stakeholders (Douali et al.,, 2022), ultimately hindering its effective implementation. Policymakers do not only have to ensure that all learners and teachers have access to these potentially powerful tools for learning - they have to provide them with basic AI literacy (Long and Magerko,, 2020; Hornberger et al.,, 2023) and the competencies they need to successfully interact with the AI systems.
5. Discussion and Conclusion
@@
Text-processing LLMs have emerged as a transformative force in the field of education ( Bhat et al., 2022; Dijkstra et al., 2023) (Abdelghani et al., 2023b, ; Ji et al.,, 2023; MacNeil et al.,, 2022). The integration of MLLMs in science educational settings holds the promise of personalized and adaptive learning experiences, capable of catering to the needs of learners by delivering content that is both accessible and engaging.
@@
MLLMs are characterized by their ability to analyze and interpret data from different sources, such as spoken and written text and images. A key advantage of MLLMs over static sources of information, such as textbook representations, is the adaptive nature of these models. Both teachers and students can use MLLMs to independently translate data, such as the evolution of ant populations, into visual representations. This capability of MLLMs can be called ”adaptive transformation.” It allows knowledge representations to alter form flexibly, especially between text and image, but also within a modality, for example, to simplify complex images or make difficult text more understandable. This shows that MLLMs can change and adapt knowledge representations almost limitlessly and with great efficiency. The question is how this can be used for knowledge acquisition.
One central question is how MLLMs and the advantages of ”adaptive transformation” can ultimately support multimedia learning, i.e., the integration of different knowledge representations into an integrated mental model. Focusing on the three stages of multimodal learning (Mayer, 2021; Schnotz & Bannert, 2003) MLLMs can play a crucial role in every stage. MLLMs and the potential of ”adaptive transformation” might help learners perform these three-step cognitive processes of selection, organization, and integration.
@@
Whereas in the past, support had to be provided by the learning material, it is now possible for learners to generate such support adaptively. For example, with MLLM it is possible for learners to summarize texts, display important text passages, or display similarities between related texts and images as key points. In addition, an adaptive use of MLLM would allow learners to integrate selection aids into multimodal representations only when they need them. This would make it possible to avoid expertise reversal (Kalyuga,, 2007) effects.
@@
A problem with employing MLLMs for (self-directed) learning might be that learners require a high degree of self-regulatory skills (e.g., (Bannert et al.,, 2014)) in order to use MLLMs for learning purposes. Learners can effectively use MLLMs to integrate multimodal learning materials. As mentioned earlier, MLLMs allow learners to select and organize learning materials according to their needs. This already prepares the way for integrating content presented in multimodal sources, as learners have better access to the learning material. In addition, multimodal representations have the potential to help learners better grasp the learning materials. For example, if learners are unable to understand complicated texts or associate complex schematic representations, MLLMs can be used to adapt learning materials so that integration is possible. It is necessary to use MLLMs in a way that the generative activities are not completely absorbed by the AI, but proportionally also by the learners. Regulating this balance seems to be a central pedagogical challenge when using MLLMs for multimedia learning.
In summary, MLLM can support learners in exploiting the potential of multimodal learning materials and presentations. In particular, through adaptive design options, either by the learners themselves or by the teachers in advance, it is possible to support learners in the selection, organization and integration of information, like text and images. This adaptivity makes learning processes potentially effective. For example, learners with a lot of prior knowledge can choose more complex representations, and learners with fewer skills can adapt learning materials to their needs. However, this also highlights the difficulty of regulating the effective use of MLLMs, either by the learners themselves or by the scripted use of AI-based applications to enable generative learning processes on the part of the learners. These issues provide an interesting area for future research and numerous potential applications.
@@
Science learning is multimodal, requiring students to engage with a variety of information formats like diagrams, tabular data, images, and animations and synthesize information as well as shift across various formats ( Unsworth, 2004) (Van Leeuwen and Kress,, 1990; Doran et al.,, 2021). A lack of representational competency can hinder science learning (Nitz et al.,, 2014). With the emergence of MLLMs, teachers now have new potential tools to address the necessity for multimodality in science education. MLLM are equipped to process and generate content beyond mere text, encompassing a spectrum of data forms, including images, audio, and video for different levels of competency. This capability is crucial in a discipline where understanding often hinges on the ability to interpret complex diagrams, analyze diverse datasets, and synthesize information from multiple sources. With MLLMs, there is potential for enhanced guidance and support in both mastering individual modalities and transitioning fluidly between them.
@@ 6. Future Implications
MLLMs could take the advantages of multimodal learning environments to the next level, as they are more flexible due to the high level of interactivity. MLLMs lack some comparability to traditional interactive multimodal learning environments because they provide an environment that is of a higher level of interactivity than typical learning environments (e.g. hypertext environment). MLLMs are able to quickly change between modes on request, something that is not possible in traditional multimedia learning environments. This could enable a highly individualized learning environment. MLLMs combine the benefits of LLMs (interactivity, real-time generation of material, etc.) with the benefits of multimodal learning (efficient use of cognitive resources). Plus, MLLMs could provide a seamless transition between modes for learners, fostering their representative competencies crucial for science understanding (Nitz et al.,, 2014).
Often, there is a shift between two modalities in the described scenarios which might be a huge shift towards more responsive, tailored to the needs of the student learning environments. As MLLMs are still in their infancy, there is a need for research. This research should not only investigate the various potentials from personalized learning to inclusiveness and more like the ones mentioned in this article. It is also crucial to focus on the potential (further) shift of the teacher’s role. As the AI systems become multimodal, they potentially become more advanced, making it relevant to think about and investigate the teacher’s role. It is crucial to understand that MLLMs, like any other AI system, should not replace the teacher or, worse, be seen as competitors, but that the MLLM alters their role in the learning process (Burbules et al.,, 2020; Schiff,, 2020). The teacher’s role changes depending on the degree of automation. Different modes and degrees of implementation of AI systems are imaginable (Molenaar,, 2022). MLLMs impact will not be limited to science education. For other disciplines, there seems to be a comparable number of opportunities and similar challenges and risks.

### v3

@@ 5. Challenges and Risks of MLLMs in Science Education
While the adaptability of MLLMs in science education is promising, it is essential to approach their potential with caution. The allure of providing students with extensive flexibility in choosing the desired modalities within their learning environment can be strong. However, evidence suggests that minimal guidance in science education may not always yield optimal learning outcomes (Kirschner et al.,, 2006; Sweller et al.,, 2007). Learners might require guidance in selecting the most appropriate modality to enhance their educational experience effectively. An abundance of options can not only serve as a distraction, but can also elevate the overall cognitive load, potentially hindering the learning process. This is especially true for students with low self-regulation capabilities (Lan,, 1998; Ley and Young,, 1998). Consequently, the roles of educators remain pivotal in guiding the learning process (Ley and Young,, 2001). They play a critical role in facilitating the effective use of these advanced tools, ensuring that the technology enhances, rather than impedes, the learning experience. In this context, the nuanced integration of MLLMs calls for a balanced approach, where the technology is leveraged thoughtfully to complement and augment traditional educational practices rather than replace them outright.
@@
Students should engage in scientific practices in problem-based learning and open inquiries (NRC,, 2012). But as minimal guidance can hinder effective learning in science education (Kirschner et al.,, 2006; Sweller et al.,, 2007) MLLMs should be designed to be adjustable by the educator to the individual student’s competency level. In this way, MLLMs would be able to provide degrees of guidance or scaffolding rather than acting as a personalized, adaptive, and multimodal agent that provides immediate answers and solutions. This mediated approach could encourage students to engage in critical thinking and problem-solving, fostering a deeper understanding and promoting self-directed learning without overexerting the student. The MLLMs should be aligned with educational and instructional strategies that prioritize the development of students’ analytical skills in tandem with their knowledge acquisition.
Ethical considerations encompass the moral and ethical implications of using MLLMs, including issues of bias, privacy, consent, and the overall impact on the quality and equity of education. In the few examples we provided, the MLLMs worked quite satisfactorily, potentially improving in the future. However, it is important to acknowledge that MLLMs sometimes still generate biased (Venkit et al.,, 2023) incorrect, faulty or fabricated content (often referred to as hallucinations) (Ji et al., 2023b, ; Zhang et al., 2023b, ; Azamfirei et al.,, 2023; Manakul et al.,, 2023), particularly in the context of logical and mathematical reasoning (Imani et al., 2023b, ). Therefore, it is crucial for researchers to empirically examine the potential bias that may contaminate assessment validity (Zhai and Nehm,, 2023). Research has already shown evidence that AI has the potential to enlarge the score divergence by gender or English language proficiency (Latif et al.,, 2023). Educators need to be aware of these limitations, as well as over-blow of AI bias (e.g., pseudo AI bias, see (Zhai and Krajcik,, 2022)). Moreover, if students are encouraged to use MLLMs independently, it is essential that educators communicate these types of potential flaws to them. Fostering a culture of ethical data handling among educational stakeholders, alongside creating awareness about the potential risks and safeguards among the student populace, is crucial to ensure a secure and responsible utilization of MLLMs in educational settings.
@@
With AI becoming an increasingly integral part of our daily lives, policymakers have begun to implement policies to regulate and safeguard AI development and its use in education (Schiff et al.,, 2020). E.g., the currently presented so-called European AI Act (European Union, (2023)) requires generative AI systems, such as MLLMs, to disclose when content is AI-generated, design models to prevent the generation of illegal content, and publish summaries of copyrighted data used for training. The use of such AI systems, and therefore MLLMs, in education will be classified as high-risk applications and will require registration in an EU database. On every level of deploying MLLMs, all stakeholders, such as developers, educators, administrators, and policymakers, as well as the corresponding governance framework, have to prioritize ethical and regulatory considerations.
@@
Educators and learners have a variety of pre- and misconceptions, fears, and hopes about AI (Bewersdorff et al., 2023b, ), which might lead to general skepticism toward AI systems in the classroom among stakeholders (Douali et al.,, 2022), ultimately hindering its effective implementation. For effective implementation, policymakers do not only have to ensure that all learners and educators have access to these potentially powerful tools for learning - they have to provide them with knowledge about AI (AI literacy: Long and Magerko, (2020); Hornberger et al., (2023)) and the competencies they need to successfully employ these AI systems.
Looking forward, there are also risks associated with proprietary MLLMs, such as unknown ethical considerations and implementations. Hence, the decision between open-source and proprietary MLLMs becomes pivotal in the long term. While proprietary models may offer robust data security and professional support, they often come with limitations in terms of accessibility, customization, and cost. This could hinder collaborative learning and innovation, as proprietary MLLMs might not be as adaptable or transparent as their open-source counterparts. Moreover, through restrictions in the ability to tailor proprietary models to specific educational contexts, such models might not always align with the diverse needs of educators, learners, researchers, or other stakeholders. As technology advances, the choice between open-source and proprietary MLLMs may not only be decisive for the trajectory of adoption of generative AI in education but also the realization of an inclusive, digitally advanced educational landscape.
6. Discussion
@@
LLMs have emerged as a transformative force in the field of education (Abdelghani et al.,, 2023; Ji et al., 2023a, ; MacNeil et al.,, 2022). The integration of MLLMs in science educational settings holds the promise of personalized and adaptive learning experiences, matching the needs of learners by delivering content that is both accessible and engaging.
@@
MLLMs are characterized by their ability to analyze and interpret data from different sources, such as spoken language, written text, or images. A key advantage of MLLMs over static sources of information, such as textbook representations, is the adaptive nature of these models. Both educators and students can use MLLMs to independently transform information into other modalities, such as textual information on the evolution of ant populations, into visual representations. This capability of MLLMs can be called ”adaptive transformation.” It allows knowledge representations to alter form flexibly, especially between text and images, but also within a modality, for example, to simplify complex images or make difficult text more understandable. This shows that MLLMs can change and adapt knowledge representations with great efficiency. One central question arising is how MLLMs can be used to foster knowledge acquisition and scientific competencies. To guide this question, we proposed a framework that is rooted in the Theory of Multimedia Learning and integrates MLLMs into multimodal learning (cf. section 3): Given visual or textual input and the user’s requirements for adaptiveness and personalization, the MLLM produces textual and/or visual outputs. The MLLM does this by transforming content from text to image or vice versa or shifting from uni- to multimodality by adding a modality. Depending on the setting, the user of the MLLM can be either the educator or the learner itself. The framework introduced in this paper proved beneficial in describing and analyzing various scenarios of applying MLLM in (science) education, including content creation, fostering student learning in diverse contexts, and providing assessment and feedback.
@@
Whereas in the past, support had to be provided by the learning material or the educator, it is now possible for learners to generate such support adaptively. For example, with MLLMs it is possible for learners to summarize texts, highlight important text passages, or display similarities between related texts and images as key points. In addition, an adaptive use of MLLMs would allow learners to integrate selection aids into multimodal representations only when they need them. This would make it possible to avoid expertise reversal effects (Kalyuga,, 2007).
@@
A problem with employing MLLMs for (self-directed) learning might be that learners require a high degree of self-regulatory skills (e.g., Bannert et al., (2014)) in order to effectively and constructively use MLLMs for learning purposes. As mentioned earlier, learners can effectively use MLLMs to integrate multimodal learning materials, as MLLMs allow learners to select, organize, and integrate learning materials according to their needs from multimodal sources. This improves the overall access to as well as the preparation of the learning material. In addition, multimodal representations have the potential to help learners better grasp the concepts to be learned. For example, if learners are unable to understand complex texts or associate complex schematic representations, MLLMs can be used to adapt learning materials so that integration is possible. It is necessary to use MLLMs in a way that the generative activities are not completely absorbed by the AI, but proportionally also by the learners. Regulating this balance seems to be a central pedagogical challenge when using MLLMs in education.
@@
Science learning is multimodal, requiring students to engage with a variety of information formats like diagrams, tabular data, natural-language texts, images, and animations and synthesize information as well as shift across various formats (Van Leeuwen and Kress,, 1990; Doran et al.,, 2021). A lack of representational competency can hinder science learning (Nitz et al.,, 2014). With the emergence of MLLMs, educators now have new potential tools to address the necessity for multimodality in science education. MLLMs are equipped to process and generate multimodal content for different levels of competency. This capability is crucial in a discipline where understanding often hinges on the ability to interpret complex diagrams, analyze diverse datasets, and synthesize information from multiple sources. With MLLMs, there is potential for enhanced guidance and support in both mastering individual modalities and transitioning fluidly between them.
In summary, MLLMs can support learners in exploiting the potential of multimodal learning materials and presentations. In particular, through adaptive design options, either by the learners themselves or by the educators in advance, it is possible to support learners in the selection, organization and integration of information contained in text, images or other modalities. This adaptivity makes learning processes potentially more effective. For example, learners with a lot of prior knowledge on a specific topic can choose more complex representations, while learners with fewer skills can adapt learning materials to their needs. However, this also highlights the difficulty of regulating the effective use of MLLMs, either by the learners themselves or by the scripted use of AI-based applications to enable generative learning processes on the part of the learners. These issues provide an area for future research and numerous potential applications.
@@ 7. Future Implications
MLLMs could take the advantages of multimodal learning environments to the next level. Unlike conventional settings, such as hypertext environments, MLLMs offer a superior degree of interactivity, enabling thus a more dynamic and engaging learning experience.
One of the most notable advantages of MLLMs is their ability to swiftly transition between different modes upon request, a feature not typically feasible in traditional multimedia learning environments. This capability paves the way for highly individualized learning experiences, tailored to the unique needs and preferences of each learner. By combining the strengths of LLMs — such as interactivity and real-time generation of content — with the efficiency of multimodal learning in optimizing cognitive resource usage, MLLMs offer new ways to support learning. Furthermore, MLLMs facilitate a seamless transition between learning modes, potentially enhancing thus the representational competencies of learners, which are crucial for a deep understanding of scientific concepts (Nitz et al.,, 2014). This seamless integration of different modalities can significantly improve the learning process.
The potential shift towards more responsive and personalized learning environments, as facilitated by MLLMs, represents a significant advancement in educational technology. This shift could lead to learning environments that are more attuned to the individual needs of students, thereby enhancing the overall learning experience.
As MLLMs are still in their infancy, there is a need for further research, which should not only investigate the various potentials from personalized learning to inclusiveness and the ones mentioned in this article. It is also crucial to focus on the potential (further) shift of the educator’s role. As AI systems become multimodal, they potentially become more advanced, making it relevant to re-think and continuously re-evaluate the educator’s role. It is crucial to understand that MLLMs, like any other AI system, should not replace the educator or, worse, be seen as competitors, but that MLLMs may alter the educators’ role in the learning process (Burbules et al.,, 2020; Schiff,, 2020). The educators’ role changes depending on the degree of automation. Different modes and degrees of implementation of AI systems are imaginable (Molenaar,, 2022). However, the impact of MLLMs will not be limited to science education; for other disciplines, there seem to be analogous opportunities and challenges (cf. Lee et al., 2023a ).
In light of these considerations, it becomes evident that the integration of MLLMs into educational frameworks across various disciplines necessitates a thoughtful and collaborative approach, where the evolving role of educators is recognized and supported, ensuring that these advanced technologies enhance, rather than overshadow, the human-centric aspect of teaching and learning.
8. Declaration of AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT (www.chat.openai.com; GPT-3.5 and GPT-4) and Grammarly (www.grammarly.com) in order to improve the readability and language of single sentences. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
9. Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.